In a blood testing procedure, blood samples from 6 people are combined into one mixture. The mixture will only test negative if all the individual samples are negative. If the probability that an individual sample tests positive is 0.12, what is the probability that the mixture will test positive?
Answer by stanbon(48568)

In a blood testing procedure, blood samples from 6 people are combined into one mixture. The mixture will only test negative if all the individual samples are negative. If the probability that an individual sample tests positive is 0.12, what is the probability that the mixture will test positive?
-------------
P(mixture tests positive) = p(at least one individual tests positive)
= 1 - P(all individuals test negative)
= 1-0.88^6
=1-0.46
=0.54
Cheers,
Stan H.